### v1

# Globally defined Carroll symmetry of gravitational waves

Abstract
The locally defined Carroll symmetry of a gravitational wave is extended to a globally defined one. Translations and Carroll boosts associated with two independent globally defined solutions of a Sturm-Liouville equation allow us to describe the motions. The Displacement Memory Effect arises for particular choices of the parameters which yield trajectories with zero momentum. We illustrate our general statements by the Pöschl-Teller profile.
pacs:
04.20.-q Classical general relativity
preprint: Global Carroll v1. arXiv: [gr-qc, hep-th]
I Introduction
Braginsky and Thorne had suggested that gravitational waves might be observed by detecting the displacement of particles initially in rest, called later the Memory Effect BraTho . While experimental verification is still in the making, an insight can be gained by taking advantage of the Carroll symmetry Leblond ; SenGupta . Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen provide a simple description Sou73 ; Carroll4GW , however suffer of being regular only in finite intervals, requiring to glue them together. In this paper we show that such complications can be avoided by switching to Brinkmann (B) coordinates Bri by solving a matrix Sturm-Liouville equation, (II.3) below. The price to pay is, though, to get more complicated-looking expressions.
We illustrate our general theory by the Pöschl-Teller profile PTeller ; Plyushchay09 ; ChaKar , which is a very good analytical approximation of the widely studied Gaussian one. We follow mostly our review EZHRev which provides the reader also with further references.
Our notations are: $({\bm{X}},U,V)$ are Brinkmann coordinates and $({\bm{x}},u,v)$ are BJR coordinates on relativistic spacetime. $U$ is an affine parameter for geodesics. $\Theta^{Brink}={\bm{\xi}}{\partial}_{{\bm{X}}}+\eta{\partial}_{V}$ denote vector fields which generate the infinitesimal isometries on spacetime.
II From Brinkmann to BJR and back
Plane gravitational waves are conveniently described by using globally defined Brinkmann coordinates Bri ; exactsol in terms of which the metric is,
$g_{\mu\nu}dX^{\mu}dX^{\nu}=d{\bm{X}}^{2}+2dUdV+K_{ij}X^{i}X^{j}\,dU^{2}\,$, (II.1)
where ${\bm{X}}=(X^{1},\dots,X^{D})$ parametrizes the transverse plane, endowed with the flat Euclidean metric $d{\bm{X}}^{2}=\delta_{ij}\,dX^{i}dX^{j}$. $U$ and $V$ are light-cone coordinates and $(K_{ij})$ we call the profile is a symmetric $D\times D$ matrix whose entries depend only on $U$. Henceforth we focus our attention at $D=1$ or $D=2$ and assume, for simplicity, that is diagonal. In $D=2$ we take
$K_{ij}X^{i}X^{j}={\mathcal{A}}(U)\Big((X^{+})^{2}-(X^{-})^{2}\Big)\,$, (II.2)
which yields a vacuum solution of the Einstein equations Bri ; exactsol . We consider “sandwich waves” BoPi89 , whose profile ${\mathcal{A}}(U)$ vanishes outside an interval $U_{i}<U<U_{f}$. The $V$-equation follows from the transverse ones EZHRev and will henceforth not be studied.
After the wave has passed, particles initially at rest exhibit the Velocity Memory Effect (VM) : they move with constant velocity. Under certain “quantization” conditions the outgoing velocity can vanish, though, and we get the Displacement Memory Effect (DM) ZelPol ; DM-1 ; DM-2 ; Sila-PLB , which might well play a rôle in future observations BraTho ; DM-2 .
Plane gravitational waves have long been known to have a $2D+1$ parameter symmetry group composed of $D+1$ translations, completed by $D$ (rather mysterious) transformations BoPiRo ; Sou73 . More recently Carroll4GW , this group was identified with the Carroll group Leblond ; SenGupta , as young Lévy-Leblond called it jokingly. Its algebraic structure is readily determined by switching to Baldwin et al (BJR) coordinates BaJe ; Rosen ; Sou73 which are however only local. This note sheds further light at the Carroll – Memory relation using globally defined Brinkmann coordinates.
We start with considering real solutions of the Sturm-Liouville equations
${P}^{\prime\prime}={\mathcal{A}}(U)P,\qquad(P^{T}){P}^{\prime}=(P^{T})^{\prime}P\,$ (II.3)
for a $D{\times}D$ matrix $P(U)=\big(P_{ij}(U)\big)$ EZHRev . The BJR coordinates $({\bm{x}},u,v)$, are then obtained by
${{\bm{X}}}=P(u){\bm{x}},\qquad U=u,\qquad V=v-\frac{1}{4}{\bm{x}}\cdot\big(P^{T}{}P\big)^{\prime}{\bm{x}}\,$, (II.4)
in terms of which the metric is,
$g_{\mu\nu}dx^{\mu}dx^{\nu}=\big(P(u)^{T}P(u)\big)_{ij}(u)\,dx^{i}dx^{j}+2du\,dv\,.$ (II.5)
Carroll symmetry consists, in BJR coordinates, of “vertical” ($v$) and transverse ($\bm{x}$}) translations with parameter $h$ and $\bf c$}, completed by Carroll boosts with parameter $\bf b$} Sou73 ; Carroll4GW ,
${\bm{x}}\to{\bm{x}}+{\bf c}+S(u)\,{\bf b}\,,\qquad u\to u,\qquad v\to v-{\bf b}\cdot{\bm{x}}-{{\scriptstyle{\frac{1}{2}}}}{\bf b}\cdot{}S(u)\,{\bf b}+h\,$, (II.6)
where the symmetric $2\times 2$ matrix
$S(u)=\int^{u}_{u_{0}}\!\left(P^{T}P\right)^{-1}\!(t)dt$ (II.7)
is referred to as the Souriau matrix Sou73 ; Carroll4GW ; EZHRev . Infinitesimal boosts parametrized with $b_{j}$ are,
$\theta^{BJR}=h\partial_{v}+c^{j}\partial_{j}+b_{j}(S^{ji}\partial_{i}-x^{j}\partial_{v})\,.$ (II.8)
Finding the trajectories is then straightforward. Noether’s theorem provides us with $2D+1$ associated conserved quantities Sou73 ; Carroll4GW ,
${\bm{p}}={\bm{p}}_{0}=(P^{T}P)(u)\,{{\bm{x}}}^{\prime}(u),\qquad{\bf k}={\bf k}_{0}={\bm{x}}(u)-S(u)\,{\bm{p}}\,$, (II.9)
interpreted as conserved linear and boost-momentum, supplemented by $m={u}^{\prime}=1$. Then the conserved quantities determine the geodesic,
${\bm{x}}(u)={\bf k}_{0}+S(u)\,{\bm{p}}_{0}\,.$ (II.10)
The subtlety comes from that the BJR coordinates are defined only in intervals $I_{k}=[u_{k-1},u_{k}]$ between the (mandatory Sou73 ; Carroll4GW ) zeros of the determinant,
$\det P(u_{k})=0\,$, (II.11)
which play a fundamentally important rôle in our investigation, as we shall see. The BJR coordinates are thus only local : both the metric (II.5) and the Souriau matrix $S$ in (II.7) are singular at the junction points $u_{k}$, as will be illustrated in FIG.1 for the Pöschl - Teller profile. The singularity problem will be resolved in sect. III by switching to Brinkmann coordinates.
Now we set $P^{T}=P$ and we consider $D=1$ (when $P$ is a scalar function) for simplicity, although our results and conclusions could be extended to non-diagonal profiles and to any dimension. Let us assume that $u<u_{1}$, the first zero of $\det(P)$. Requiring that our particles be in rest before the wave arrives,
${{\bm{x}}}(-\infty)=x_{0}{\quad\text{\small and}\quad}{{\bm{x}}}^{\prime}(-\infty)=0\,$, (II.12)
imply that the momentum vanishes in $I_{1}$,
${\bm{p}}=0\,.$ (II.13)
The Souriau term is thus switched off from (II.10) and the transverse BJR “trajectory” is merely a fixed point.
${\bm{x}}(u)={\bm{x}}_{0}=\mathop{\rm const}\nolimits\,$, (II.14)
confirming, in BJR coordinates, the “no motion for Carroll” maxim Leblond . In contrast, the Brinkmann trajectory is non-trivial : (II.4) yields,
${\bm{X}}(U)=P(u)\,{\bm{x}}_{0}\,$, (II.15)
with $U=u$. The initial conditions (II.12) then require :
$P(-\infty)=1{\quad\text{\small and}\quad}P^{\prime}(-\infty)=0\,.$ (II.16)
The Brinkmann trajectory, given by the $P$-matrix, may look quite complicated, depending on the Sturm-Liouville solution $P$ of (II.4) EZHRev . For appropriate “quantized” values of the wave parameter the wave zone contains an integer number of half-waves; then $P(u)$ tends to a constant matrix also for $u\to\infty$ : DM is obtained DM-1 ; DM-2 ; Sila-PLB .
We emphasise that the investigations above are a priori valid only before the first zero of $\det{P}$ and should then be restarted until the next zero, and so on.
Another remarkable consequence of (II.15) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from an $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin, and we get a caustic point BoPi89 ; EZHRev . The question will be further discussed in Sec.IV and illustrated in FIG. 2.
III Carroll symmetry in Brinkmann coordinates
Simple as they are, BJR coordinates (II.4) suffer from being valid only in intervals between subsequent zeros of $P$, and the adjacent expressions should be glued together – which may be laborious. A global approach can conveniently be given instead by using Brinkmann coordinates. For simplicity we restrict our attention at $D=1$ transverse dimension and start with a fixed interval $I_{k}$. A straightforward computation then shows that pulling back the BJR expression (II.6) by using (II.4) [or its infinitesimal form (II.8)] yields the Carroll generators written in Brinkmann coordinates,
${\Theta^{Brink}=h\frac{\partial}{\partial V}+c\left(P\frac{\partial}{\partial X}-P^{\prime}X\frac{\partial}{\partial V}\right)+b\left((PS)\frac{\partial}{\partial X}-(PS)^{\prime}X\frac{\partial}{\partial V}\right)\,,}\;\;$ (III.1)
The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the the symmetry generators $\xi$ and $\eta$, mentioned in the Introduction. In Sec.IV they will be spelt out explicitly for Pöschl - Teller .
The $P$-matrix and its derivative $P^{\prime}$ generate translations, whereas $PS$ and $(PS)^{\prime}$ generate Carroll boosts GenRelGrav . Their only nonzero commutator yields a vertical translation Blau24 ,
$\Big[\underbrace{P\frac{\partial}{\partial X}-P^{\prime}X\frac{\partial}{\partial V}}_{translation}\ ,\ \underbrace{PS\frac{\partial}{\partial X}-(PS)^{\prime}X\frac{\partial}{\partial V}}_{boost}\Big]=-\partial_{V}\,.$ (III.2)
The symmetry is thus the Heisenberg algebra.
$P$ thus plays a double role : by (II.15) it determines the B trajectory, whereas (III.1) says that it also generates translations. The B trajectory is obtained indeed from $x_{0}=\mathop{\rm const}\nolimits$ by a (generally $U$-dependent) translation.
The Brinkmann form (III.1) has an important advantage w.r.t. the BJR expression in (II.6) : multiplication by $P$ supresses the singularity of Souriau matrix $S$ Sou73 ; EZHRev  :
Q(U)=P(U)S(U)Q(U)=P(U)S(U)\, (III.3)
is regular for all $U$. Thus (III.1) extends the Carroll symmetry (II.6) naturally from a fixed interval $I_{k}$ to the entire $U$-axis.
Both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.3) EZHRev ; Blau24 as it can be verified by a direct calculation. The Wronskian is $W(P,Q)=P^{\prime}Q-Q^{\prime}P=-1$, therefore $P$ and $Q$ are independent. The expression (III.1) could also be verified independently, by checking the symmetry equation $L_{Y}g=0$ Torre ; GenRelGrav .
Contracting the Brinkmann metric (II.1) with (III.1) provides us with,
${\bm{p}}_{0}=P{\bm{X}}^{\prime}-\big(P\big)^{\prime}{\bm{X}}{\quad\text{\small and}\quad}{\bf k}_{0}=-Q{\bm{X}}^{\prime}+\big(Q\big)^{\prime}{\bm{X}}\,$, (III.4)
where ${\bm{p}}_{0}$ and ${\bf k}_{0}$ are, a priori, new constants of the motion. However expressing them in BJR terms, (II.4), the previous expressions (II.9) are recovered; as anticipated by our notation.
The Sturm-Liouville solutions $P$ and $Q$ play also yet another rôle, though : either combining (II.10) with (II.4) or solving (III.4) directly yields the general Brinkmann trajectory,
${{\bm{X}}(U)=P(U)\,{\bf k}_{0}+Q\,{\bm{p}}_{0}\,.}\;\;$ (III.5)
which is our second main result. When ${\bm{p}}_{0}=0$ cf. (II.13), then the $Q$ - term is switched off and we recover (II.15) with ${\bm{x}}_{0}={\bf k}_{0}$. For parameters which correspond to having an integer number of half-waves in the wavezone, we have also $P(+\infty)=\mathop{\rm const}\nolimits$ and we get DM.
IV Illustration by the Pöschl - Teller profile
We illustrate our general theory by the Pöschl - Teller profile PTeller ; Plyushchay09 ; ChaKar ; DM-1 in $D=1$,
${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{4m(m+1)}{\cosh^{2}U}\,.$ (IV.1)
The Sturm-Liouville equation (II.3) is solved by Legendre functions. DM trajectories arise when $m$ is an integer, and we get Legendre polynomials, $(-1)^{m}P_{m}(\tanh U)$ DM-1 which have $m$ zeros.
Turning to BJR by (II.4), the Souriau matrix $S$ in (II.7) is regular between two subsequent zeros of $P$ but diverges at the junction points. For $m=1$, for example,
$S_{m=1}(u)=u-\coth u\,$ (IV.2)
obtained for the choice $u^{\pm}_{0}\approx\pm 1.2$, is regular either in $I_{-}<0$ or in $I_{-}>0$, but diverges at $u=0$, as depicted in FIG.1a. Similarly for $m=2$,
$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$, (IV.3)
is singular where the denominator vanishes, as shown in FIG.1b.
Figure 1: The Sturm-Liouville solutions $\bf P$ in (II.3) are DM trajectories written in Brinkmann coordinates with wave numbers $\bf m=1$} and $\bf m=2$}, respectively. The Souriau matrix S diverges where $\bf P=0$, but the pull-back to Brinkmann $\color[rgb]{0,0.5,0}\bf Q=PS$} is regular for all $U$.
A second solution of the Sturm-Liouville equation (II.3), $Q_{m}=P_{m}S_{m}$ , is regular for all $U$ but it is a Legendre function with non-DM behavior. The solutions $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$. For $m=1$ and $2$ we have, for example,
$\begin{array}[]{llclllc}P_{1}(U)&=&{-}\tanh U\,,&\qquad&P_{2}(U)&=&\frac{1}{2}(3\tanh^{2}U-1)\,,\\[6.0pt] Q_{1}(U)&=&{1-U\tanh U},&\qquad&Q_{2}(U)&=&\frac{1}{2}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\,.\end{array}$ (IV.4)
The two SturmLiouville{Sturm-Liouville\;} solutions are shown in FIG. 1 in red and in green.
The $P(U)$ found above provide us with DM geodesics which sare consistent with (II.15). Their caustic behavior at the zeros of $P$ is manifest in FIG.2.
Figure 2: For a DM wave in $D=1$ transverse dimension Brinkmann trajectories focus to the caustic points determined by $P_{m}(U_{k})=0$, $k=1,\dots,m$. For $m=1$ there is one focal point at the origin, and for $m=2$, we have two of them, determined by $\tanh U_{\pm}=\pm{1}/{\sqrt{3}}$.
Coming to symmetries of Pöschl - Teller , the coefficients of the Carroll generators $P$ and $Q$ in (III.1) are shown in FIGs. 3 and 4.
(a)                       (b)
Figure 3: For wave number $\bf m=1$}, the (a) translation generator ${\bf P}_{1}$ [which is also a trajectory] is odd. (b) the generator of a Carroll boost, ${\bf Q}_{1}={\bf P_{1}S_{1}}$, is even. The only zero of ${\bf P}_{1}$ is at $U=0$.
(a)                       (b)
Figure 4: Infinitesmal Carroll generators (IV) with wave number $\bf m=2$}. (a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$. (b) For a boost, ${\bf Q}_{2}={\bf P_{2}S_{2}}$ is odd.
The globally defined Carroll symmetry generators (III.1) are
$\Theta_{m=1}^{Brink}$ $=$ $h\frac{\partial}{\partial V}+c\Big({\tanh U}\,\frac{\partial}{\partial X}-(\operatorname{sech}^{2}U)X\frac{\partial}{\partial V}\Big)$ (IV.5)
$+$ $b\Big({\big(U\tanh U-1\big)}\frac{\partial}{\partial X}-\big(\tanh U+U\operatorname{sech}^{2}U\big)X\frac{\partial}{\partial V}\Big)\,.$
$\Theta_{m=2}^{Brink}\!$ $=$ $\!h\frac{\partial}{\partial V}+c\left(\big(3\tanh^{2}U-1\big)\frac{\partial}{\partial X}-\big(6\tanh U\operatorname{sech}^{2}(U)\big)X\frac{\partial}{\partial V}\right)$
$\thinspace{+}b\left(\frac{1}{4}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\frac{\partial}{\partial X}-\frac{1}{2}\big(1-3\operatorname{sech}^{2}U+3U\operatorname{sech}^{2}U\tanh U\big)X\frac{\partial}{\partial V}\right).\qquad$
V Planar representation of the symmetry generators
The symmetry-generating vectors along the trajectories, shown in FIGs.3 and 4, are conveniently viewed in a co-moving tangent space – which is indeed a plane carried along the trajectory. In FIGs. 58 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ of the symmetry generators in (III.1), $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$, spelt out in (IV.5)-(IV).
Figure 5: Infinitesimal translations $({\bm{\xi}},\eta)$ along a DM Brinkmann trajectory ${\bm{X}}(U)$ with $\bf m=1$}, shown in the co-moving tangent space. The trajectory is “hidden” in the yellow blob.
FIG.5 shows that the translation vector starts at $\bf U=-\infty$} from $\color[rgb]{.75,.5,.25}{\bm{\xi}}=-1,\eta=0$}. It leaves the caustic point for $\bf U=0$} at ${\bm{X}}=0,\,V=0$ invariant. Arriving into the Afterzone, the translation vectors change sign and end, for $U=+\infty$, at $\color[rgb]{0,1,1}{\bm{\xi}}=+1,{\eta}=0$}. Thus they act again as usual translations – but with reversed sign.
Figure 6: Infinitesimal boosts along along the trajectory “hidden behind the yellow blob” with $\bf m=1$}, shown in the co-moving tangent space.
In FIG.6, the longest arrow (in brown) is both the initial and the final boost vector for $\bf U=\mp\infty$}. The shortest red arrow $({\bm{\xi}}=1,{\eta}=0)$ is reached for $\bf U=0$} at the caustic point ${\bm{X}}=0,V=0$, where the boost shifts all trajectories by the same amount.
Figure 7: Translations act along the $m=2$ DM trajectory by (IV) The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
According to FIG.7, the translation generator (with $c=1/2$) is, for $U=$-$\infty$ , $({\bm{\xi}}=1,\,\eta=0)$. FIGs. 3 and 4 confirm that, consistently with (IV.5), the two focal points of the trajectory, at $U_{\mp}$, are left invariant. For $U=+\infty$ the “8-shaped” curve returns to $({\bm{\xi}}=1,\,\eta=0)$, where it started from.
Figure 8: Infinitesimal Carroll boosts in Brinkmann coordinates along a chosen DM trajectory marked by a yellow blob, with wave number $\bf m=2$}.
Boosts along the $m=2$ trajectory are shown in. FIG.8. The longest vector on the left (in brown) is, consistently with (IV.5) and (IV) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$. The two shortest blue arrows (${\bm{\xi}}=\pm 1,{\eta}=0$) show the boost acting as a translations by $\mp b$ at the caustic points $U_{\mp}$.
VI Conclusion
Our results shed further light on the relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.16) yields our geodesics. The general expression (III.5) which involves both DM and VM is one of our main results.
When the parameter $m$ is an integer, then the wavezone accommodates $m$ half-waves and we get DM. For $m=2\ell+1$ $P(U)$ is odd and we do get net displacement. For $m=2\ell$, though, $P(U)$ is even and the particle returns, after non-trivial motion in the wave zone, to its initial position : we get DM with no final displacement, consistently with (II.15) and confirmed by FIGs 2, 3a and 4a.
DM trajectories are distinguished by having zero momentum and can be found by putting ${\bm{p}}_{0}=0$ into (III.5). Turning on the momentum, ${\bm{p}}_{0}\neq 0$, distorts the trajectory to VM.
The merit of BJR coordinates is their simplicity Sou73 ; Carroll4GW . For our initial conditions (II.16), the “motion” reduces to a fixed point, given by the conserved boost momentum ${\bm{x}}_{0}={\bf k}$ in (II.10). The entire dynamics is carried by $P(u)$. The price to pay is that the BJR description works only between zeros of the $P$-matrix, and must then be fitted together. These same points, distinguished by the vanishing of $\det(P)$.
The advantage of the Brinkmann form (III.1), which is one of our principal results, is that it replaces the necessarily singular Souriau matrix $S$ Sou73 by the globally defined Sturm-Liouville solution $Q$. As examplified by the Pöschl - Teller profile, it allowed us to extend the locally given Carrollian Killing vectors to entire spacetime. The clue is that the multiplying by $P$ cures the singularity of the Souriau matrix $S$ and thus that of the BJR coordinate system. The Brinkmann trajectories are focused at the zeros of $\det(P)$, as shown in FIG.2.
The Carroll symmetry generators, first identified by Souriau Sou73 , look more complicated in Brinkmann coordinates as their BJR counterparts do. They are however defined globally. The are directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by using the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.5). The intimate relationship of geodesics with Carroll symmetry is highlighted by their common use of the Sturm-Liouville equation GenRelGrav .
Our results go actually beyond $D=1$ dimensional profiles and the Pöschl - Teller example. Similar results hold for the Scarf profile in $D=2$, which yields a good analytic approximation of flyby ZelPol ; Sila-PLB .
The supersymmetric extension MSUSY sheds further light on the Memory Effect. Further details will be presented in a forthcoming comprehensive review ZEHPR .
Acknowledgements.
The authors are indebted to J. Balog for correspondance. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320). ME is supported by Istanbul Bilgi University research fund (BAP) with grant no: 2024.01.009 and by The Scientific and Technological Research Council of Turkey (TÜBITAK) under grant number 125F021. PMZ and PAH benefitted of TÜBITAK 2221 grant and hospitality at Bilgi University during their visit in Istanbul.