### v3

# Globally defined Carroll symmetry of gravitational waves

Abstract
The local Carroll symmetry of a gravitational wave found in Baldwin-Jeffery-Rosen coordinates is extended to a globally defined one by switching to Brinkmann coordinates. Two independent globally defined solutions of a Sturm-Liouville equation allow us to describe both the symmetries (translations and Carroll boosts) and the geodesic motions. One of them satisfies particular initial conditions which imply zero initial momentum, while the other does not. Pure displacement arises when the latter is turned off by requiring the momentum to vanish and when the wave parameters take, in addition, some particular values which correspond to having an integer half-wave number. The relation to the Schwarzian derivative is highlighted. We illustrate our general statements by the Pöschl-Teller profile.
pacs:
04.20.-q Classical general relativity
I Introduction
Braginsky and Thorne BraTho argued that gravitational waves might be observed by detecting the displacement of particles initially in rest, called later the Memory Effect. While experimental verification is still in the making, conceptual insight can be gained by taking advantage of the Carroll symmetry LeblondCar ; SenGupta . Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen provide a simple description Sou73 ; Carroll4GW , however are regular only in finite intervals, requiring to glue together the local results.
In this paper we show that such complications can be avoided by switching to globally defined Brinkmann (B) coordinates Bri ; DGH91 . The transcription requires to solve a matrix Sturm-Liouville equation, (II.6) below. The price to pay for globality is to get more complicated-looking expressions, though. The advantages and drawbacks of the two coordinate systems is discussed in the Conclusion.
We illustrate our general theory by the Pöschl-Teller profile PTeller ; Plyushchay09 ; ChaKar , which is a good analytical approximation of the widely studied Gaussian one. The book exactsol and our review EZHRev provide the reader with further references.
Our notations are: $({\bm{X}},U,V)$ are Brinkmann coordinates and $({\bm{x}},u,v)$ are BJR coordinates. $U=u$ is an affine parameter for geodesics. $\Theta^{Brink}={\bm{\xi}}{\partial}_{{\bm{X}}}+\eta{\partial}_{V}$ denotes collectively the vector fields which generate the infinitesimal isometries of spacetime. Its explicit form will be spelt out in (III.1).
II From Brinkmann to BJR and back
Plane gravitational waves are conveniently described by using globally defined Brinkmann coordinates Bri ; DGH91 ; exactsol in terms of which the metric is,
$g_{\mu\nu}dX^{\mu}dX^{\nu}=d{\bm{X}}^{2}+2dUdV+K_{ij}X^{i}X^{j}\,dU^{2}\,$, (II.1)
where the ${\bm{X}}=(X^{1},\dots,X^{D})$ are coordinates of the transverse plane with the flat Euclidean metric $d{\bm{X}}^{2}=\delta_{ij}\,dX^{i}dX^{j}$. $U$ and $V$ are light-cone coordinates related to the usual relativistic coordinates according to $U=\frac{1}{\sqrt{2}}(Z-T)$ and $V=\frac{1}{\sqrt{2}}(Z+T)$. The profile $(K_{ij})$ is a symmetric $D\times D$ matrix whose entries depend only on $U$. Henceforth we focus our attention at $D=1$ or $D=2$ and assume, for simplicity, that the wave is linearly polarized so that $(K_{ij})$ is diagonal. Thus in $D=2$, which is our main interest in this paper,
$K_{ij}X^{i}X^{j}={\mathcal{A}}(U)\Big((X^{+})^{2}-(X^{-})^{2}\Big)\,$, (II.2)
which yields a vacuum solution of the Einstein equations for any ${\mathcal{A}}(U)$ Bri ; exactsol ; DGH91 ; EZHRev . We consider only “sandwich waves” BoPi89 , whose profile ${\mathcal{A}}(U)$ vanishes outside an interval $U_{i}<U<U_{f}$. Choosing $U$ as an affine parameter, the equations of motion of the transverse coordinates are of the Sturm-Liouville form,
$\dfrac{d^{2}{\bm{X}}}{dU^{2}}-\left(\begin{array}[]{lr}{\mathcal{A}}&0\\ 0&-{\mathcal{A}}\end{array}\right)\,{\bm{X}}=0\,.$ (II.3)
The geodesic equations have an additional constant of the motion called the Jacobi invariant exactsol ,
$e=\frac{1}{2}\,g_{\mu\nu}\,\frac{d{X}^{\mu}}{dU}\frac{d{X}^{\nu}}{dU}\,.$ (II.4)
The transverse equations (II.3) do not depend on $e$. In this paper we limit our attention at the lightlike case $e=0$ for which the “vertical” coordinate $V$ satisfies
$\dfrac{d^{2}V}{dU^{2}}+\dfrac{1}{2}\displaystyle{\frac{d{\mathcal{A}}}{dU\,}}\Big((X^{1})^{2}-(X^{2})^{2}\Big)+2{\mathcal{A}}\Big(X^{1}\dfrac{dX^{1}}{dU\,}-X^{2}\dfrac{dX^{2}}{dU\,}\Big)=0\,.$ (II.5)
It corresponds to the horizontal lift of (II.3) and can be solved once the solutions of the latter are known. Henceforth we focus our attention at (II.3). The timelike (massive) case, $e<0$, is discussed, e.g. in EDAHK ; DM-1 . The equations above constitute a model for particle dynamics in a gravitational wave including flyby ZelPol .
After the wave has passed, particles initially at rest exhibit the Velocity Memory Effect (VM) Ehlers ; AiBalasin ; EZHRev  : outside the Wavezone, they move with constant velocity. Under certain “quantization” conditions the outgoing velocity can vanish, though, and we get the Displacement Memory Effect (DM) ZelPol ; DM-1 ; DM-2 ; Sila-PLB .
Plane gravitational waves have long been known to have a $2D+1$ parameter symmetry group composed of $D+1$ translations, completed by $D$ (rather mysterious) transformations BoPiRo ; Sou73 . More recently Carroll4GW , this group was identified with the Carroll group LeblondCar ; SenGupta , as young Lévy-Leblond called it jokingly. Its algebraic structure is readily determined by switching to Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen ; Sou73 which are however only local. This note sheds further light at the Carroll — Memory relation using globally defined Brinkmann coordinates.
We start with considering real solutions of the Sturm-Liouville equations
${P}^{\prime\prime}={\mathcal{A}}(U)P,\qquad(P^{T}){\,}{P}^{\prime}=(P^{T})^{\prime}{\,}P\,$ (II.6)
for a $D{\times}D$ matrix $P(U)=\big(P_{ij}(U)\big)$ EZHRev . The prime here denotes $d/du\equiv d/dU$. BJR coordinates $({\bm{x}},u,v)$ are then obtained by
${{\bm{X}}}=P(u){\bm{x}},\qquad U=u,\qquad V=v-\frac{1}{4}{\bm{x}}\cdot\big(P^{T}{}P\big)^{\prime}{\bm{x}}\,$, (II.7)
in terms of which the metric takes the BJR form,
$g_{\mu\nu}dx^{\mu}dx^{\nu}=\left(P(u)^{T}P(u)\right)_{ij}(u)\,dx^{i}dx^{j}+2du\,dv\,.$ (II.8)
Carroll symmetry written in BJR coordinates is composed of “vertical” ($v$) and transverse ($\bm{x}$}) translations with parameters $h$ and $\bf c$}, respectively, completed by Carroll boosts with parameter $\bf b$} Sou73 ; Carroll4GW ,
${\bm{x}}\to{\bm{x}}+{\bf c}+S(u)\,{\bf b}\,,\qquad u\to u,\qquad v\to v-{\bf b}\cdot{\bm{x}}-{{\scriptstyle{\frac{1}{2}}}}{\bf b}\cdot{}S(u)\,{\bf b}+h\,.$ (II.9)
The symmetric $2\times 2$ matrix
$S(u)=\int^{u}_{u_{0}}\!\left(P^{T}P\right)^{-1}\!(t)dt$ (II.10)
here is referred to as the Souriau matrix Sou73 ; Carroll4GW ; EZHRev . $S$ is related to but is slightly different from the time redefinition considered, in the isotropic case, by Arnold Arnold ; ZurabNote ; Aniso ; ZZH . Details will be as discussed elsewhere ZEHPR .
The infinitesimal symmetries are, with some abuse of notation,
$\theta^{BJR}=h\partial_{v}+c^{j}\partial_{j}+b_{j}\big(S^{ji}\partial_{i}-x^{j}\partial_{v}\big)\,.$ (II.11)
The covariantly constant Killing vector ${\partial}_{v}$ here generates “vertical” translations – a hallmark of Brinkmann metrics Bri ; DGH91 . The $c_{i}$ generate translations and the $b_{j}$ generate Carroll boosts. For $u=u_{0}$, the integral (II.10) vanishes and we recover the usual Carroll action in LeblondCar .
Finding the geodesic trajectories is then straightforward by using the symmetry. Noether’s theorem provides us with associated conserved quantities Sou73 ; Carroll4GW ,
${\bm{p}}={\bm{p}}_{0}=(P^{T}P)(u)\,{{\bm{x}}}^{\prime}(u),\qquad{\bf k}={\bf k}_{0}={\bm{x}}(u)-S(u)\,{\bm{p}}\,$, (II.12)
interpreted as conserved linear and boost-momentum, respectively, supplemented by the non-relativistic mass that we scale to $1$. It is generated by the Killing vector ${\partial}_{v}$. The conserved quantities determine the geodesic trajctories  [^4] A non-vanishing Jacobi invariant $e\neq 0$ would merely add a linear term, $eu$ to the r.h.s., EDAHK ; DM-1 .,
${\bm{x}}(u)={\bf k}_{0}+S(u)\,{\bm{p}}_{0}\,,\quad v(u)=-\frac{1}{2}{\bm{x}}.{\bm{p}}+v_{0}\,.$ (II.13)
So far so good. However a subtlety comes from that the BJR coordinates are defined only in intervals $I_{k}=[u_{k-1},u_{k}]$ between mandatory Sou73 ; Carroll4GW zeros of the determinant [^5] In $D=1$, $P(U)$ is just a scalar function.,
$\det P(u_{k})=0\,.$ (II.14)
The matrix $P$ plays a fundamentally important rôle in our investigation, as we shall see. Both the metric (II.8) and the Souriau matrix $S$ (II.10) are singular at the junction points $u_{k}$, as is manifest in FIG.2 for the Pöschl - Teller profile. The singularity will be removed by switching to Brinkmann coordinates.
Although our results could be extended to non-diagonal profiles and to any dimension, henceforth we focus our attention at the diagonal case, $P^{T}=P$, and to D=2D=2i.e.,  to the physically relevant $4$-dimensional spacetime. Occasionally, we consider also toy examples in $D=1$, though.
Let us assume that $u<u_{1}$, the first zero of $\det(P)$. Requiring that our particles be in rest before the wave arrives,
${{\bm{x}}}(-\infty)=x_{0}{\quad\text{\small and}\quad}{{\bm{x}}}^{\prime}(-\infty)=0\,$, (II.15)
implies that the momentum vanishes in $I_{1}$,
${\bm{p}}=0\,.$ (II.16)
The Souriau term is thus switched off from (II.13) and the transverse BJR “trajectory” is, for vanishing Jacobi invariant, $e=0$ in (II.4), merely a fixed point  [^6] A non-vanishing Jacobi in $e\neq 0$ would merely shift $v(u)$ by a linear-in-$u$ term EDAHK ; DM-1 .,
${\bm{x}}(u)={\bm{x}}_{0}=\mathop{\rm const}\nolimits\,,\qquad v(u)=v_{0}=\mathop{\rm const}\nolimits\,$, (II.17)
consistently with (II.13). Eqn (II.17) confirms the “no motion for Carroll” wisdom LeblondCar   [^7] Motion is possible for multi-particle systems MultiCarroll , and for planar models with “exotic” double central extension when the particles follow the Hall law MARSOT ; ExoPeierls .. However, unlike as for BJR, the (lightlike) Brinkmann trajectory is non-trivial : (II.7) yields,
${\bm{X}}(U)=P(u)\,{\bm{x}}_{0}\,,\qquad V(U)=-\frac{1}{2}{\bm{X}}.{\bm{X}}^{\prime}+V_{0}\,$, (II.18)
with $U=u$ and $V_{0}=v_{0}$. The initial conditions (II.15) then require :
$P(-\infty)={\rm Id}{\quad\text{\small and}\quad}P^{\prime}(-\infty)=0\,.$ (II.19)
The Brinkmann trajectory, given by the action of the $P$-matrix, (II.18), may look indeed quite complicated, depending on the Sturm-Liouville solution $P$ of (II.6) EZHRev . For randomly chosen wave parameters we get VM Ehlers ; AiBalasin ; EZHRev  : the particles fly apart with constant velocity. For appropriately “quantized” values of the wave parameter, though, the wave zone contains an integer number of half-waves; $P(u)$ then tends to a constant matrix also for $u=U\to\infty$ and DM is obtained DM-1 ; DM-2 ; Sila-PLB .
We emphasise, though, that the investigations above are a priori valid only before the first zero of $\det{P}$ and should then be restarted until the next zero, and so on, as it will be seen in sec.V for Pöschl - Teller .
Another remarkable consequence of (II.18) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin : we get a caustic point BoPi89 ; EZHRev , as it will be illustrated in FIG. 3.
III Carroll symmetry in Brinkmann coordinates
Simple as they are, BJR coordinates (II.7) suffer from being valid only in intervals between subsequent zeros of $det(P)$. The adjacent expressions should be glued together – which may be laborious. A global approach can conveniently be given instead by using Brinkmann coordinates. For simplicity we restrict our attention at $D=1$ transverse dimension and start with a fixed interval $I_{k}$. A straightforward computation then shows that pulling back the BJR expression (II.9) by (II.7) [or its infinitesimal form (II.11)] yields the Carroll generators written in Brinkmann coordinates,
${\Theta^{Brink}=h\frac{\partial}{\partial V}+c\left(P\frac{\partial}{\partial X}-P^{\prime}X\frac{\partial}{\partial V}\right)+b\left((PS)\frac{\partial}{\partial X}-(PS)^{\prime}X\frac{\partial}{\partial V}\right)\,,}\;\;$ (III.1)
which could also be verified independently, by checking the equation $L_{Y}g=0$ Torre ; GenRelGrav . The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the symmetry generators $\xi$ and $\eta$ mentioned in the Introduction. They will be spelt out explicitly for Pöschl - Teller in Sec.V. The $P$-matrix and its derivative $P^{\prime}$ generate translations, whereas $Q=PS$ and $Q^{\prime}=(PS)^{\prime}$ generate Carroll boosts GenRelGrav . ${\partial}_{V}$ generates the non-relativistic mass scaled to $1$, as does for BJR.
The only nonzero commutator yields a vertical translation Blau24 ,
$\Big[\underbrace{P\frac{\partial}{\partial X}-P^{\prime}X\frac{\partial}{\partial V}}_{translation}\ ,\ \underbrace{PS\frac{\partial}{\partial X}-(PS)^{\prime}X\frac{\partial}{\partial V}}_{boost}\Big]=-\partial_{V}\,.$ (III.2)
The symmetry is thus the Heisenberg algebra.
$P(U)$ thus plays a double role : for DM parameters it determines the complicated-looking Brinkmann trajectory in (II.18), whereas by (III.1) it also generates translations. For non-DM parameters, the situation is even more complicated, as will be seen in (III.6).
The Brinkmann form (III.1) has an important advantage w.r.t. the BJR expression (II.9) : left multiplication by $P$ removes the singularity of the Souriau matrix $S$ Sou73 ; EZHRev  :
Q(U)=P(U)S(U)Q(U)=P(U)S(U)\, (III.3)
is regular for all $U$. In conclusion, switching form BJR to Brinkmann, (III.1),extends the Carroll symmetry (II.9) from an interval $I_{k}$ to the entire $U$-axis. A remarkable observation says that both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.6),
${Q}^{\prime\prime}={\mathcal{A}}(U)Q,\qquad(Q^{T}){\,}{Q}^{\prime}=(Q^{T})^{\prime}{\,}Q\,$, (III.4)
as it can also be verified by a direct calculation. $P$ and $Q$ have different initial conditions, though EZHRev ; Jibril ; Blau24 . The two solutions are independent when their Wronskian, $W(P,Q)=P^{\prime}Q-Q^{\prime}P$, does not vanish.
Contracting the Brinkmann metric (II.1) with the Carroll generators (III.1) provides us with,
${\bm{p}}_{0}=P{\bm{X}}^{\prime}-\big(P\big)^{\prime}{\bm{X}}{\quad\text{\small and}\quad}{\bf k}_{0}=-Q{\bm{X}}^{\prime}+\big(Q\big)^{\prime}{\bm{X}}\,$, (III.5)
where ${\bm{p}}_{0}$ and ${\bf k}_{0}$ are, a priori, new constants of the motion. However expressing them in BJR terms using (II.7), the previous expressions (II.12) are recovered, as anticipated by our notation.
The Sturm-Liouville solutions $P$ and $Q$ play yet another rôle, though : either combining (II.13) with (II.7) or solving (III.5) directly yields the general Brinkmann trajectory,
${{\bm{X}}(U)=P(U)\,{\bf k}_{0}+Q(U)\,{\bm{p}}_{0}\,.}\;\;$ (III.6)
Remember that $P$ satisfies the initial conditions (II.19).
For vanishing momentum ${\bm{p}}_{0}=0$, cf. (II.16), the VM term $Q$ is switched off and we recover (II.18) with ${\bm{x}}_{0}={\bf k}_{0}$. The corresponding Brinkmann trajectory may or may not be DM yet, though, depending on the wave profile. For parameters for which we have an integer number of half-waves in the wavezone, we do get DM : $P(+\infty)=\mathop{\rm const}\nolimits$, though DM-1 ; DM-2 ; Sila-PLB . The case of nonzero initial momentum, ${\bm{p}}_{0}\neq 0$, will be studied in ZEHPR .
IV Schwarzian aspects
For simplicity, we work in $D=1$. Our results have an interesting relation with the Schwarzian derivative defined for a function $f$ as Ovsienko ; Vlahakis ; ZZH ,
${\mathfrak{S}}\big(f(x)\big)=\frac{f^{\prime\prime\prime}(x)}{f^{\prime}(x)}-\frac{3}{2}\left(\frac{f^{\prime\prime}(x)}{f^{\prime}(x)}\right)^{2}\,.$ (IV.1)
We note for later use that the Schwarzian depends only on $f^{\prime}$, not on $f$ itself.
Consider now the Sturm-Liouville equation
$\varphi^{\prime\prime}(x)-{\mathcal{A}}(x)\varphi=0\,$, (IV.2)
cf. (II.6) for both $P$ or for $Q$ collectively. Let then $\varphi_{1}(x)$ and $\varphi_{2}(x)$ be two independent solutions and consider their quotient,
$f(x)=\frac{\varphi_{1}(x)}{\varphi_{2}(x)}\,.$ (IV.3)
It is then shown in Ovsienko that the Sturm-Liouville profile can be recovered from the solutions in terms of the Schwarzian derivative,
${{\mathcal{A}}=-{\scriptstyle{\frac{1}{2}}}\,{\mathfrak{S}}(f)\,.}\;\;$ (IV.4)
Let us now apply these general formulas with the cast $x\leadsto U$, completed by,
$\varphi_{1}(U)=P(U){\quad\text{\small and}\quad}\varphi_{2}(U)=P(U)S(U)\,.$ (IV.5)
Thus
$f=\frac{1}{S}\,.$ (IV.6)
The Wronskian $W(P,Q)$ is a constant, implying that $f^{\prime}$ in (IV.1) is indeed proportional to $S^{-2}.$ Thus by (IV.4) the profile is the Schwarzian derivative of the inverse of the Souriau matrix,
${\mathcal{A}}=-\frac{1}{2}\,{\mathfrak{S}}\left(\frac{1}{S}\right)\,.$ (IV.7)
The form (II.10) of $S$ then implies that $P$ determines also the second solution we called $Q$.
Remarkably, we can also interchange the cast: choosing
$\varphi_{1}(U)=Q=P(U)S(U){\quad\text{\small and}\quad}\varphi_{2}(U)=P(U)\,$ (IV.8)
instead of (IV.5), it is $P$ which drops out, leaving us with the reciprocal relation for
$\tilde{f}=\frac{Q}{P}=S\,$, (IV.9)
instead of $f=P/Q=1/S$ as in (IV.6).
The new cast (IV.8) then yields, once again, a SL eqn. (IV.2) whose profile is determined by the Souriau matrix,
${\mathcal{A}}(U)=-2{\mathfrak{S}}(S)=-2\frac{S^{\prime\prime\prime}(U)}{S^{\prime}(U)}+{3}\left(\frac{S^{\prime\prime}(U)}{S^{\prime(}U)}\right)^{2}\,.$ (IV.10)
which can indeed be viewed as an equation for $S$. Somewhat surprisingly, inserting here $S$ from (II.10), the previous Sturm-Liouville equation (II.6) with the same profile $\mathcal{A}$} is recovered. These results indicate that the interchange
$P,\,Q\to Q,\,P\;\;\;\Rightarrow\;\;\;S\to\frac{1}{S}$ (IV.11)
is a symmetry of the system, as it follows from the reciprocity relation
${\mathfrak{S}}(f)={\mathfrak{S}}\left(\frac{1}{f}\right)$ (IV.12)
of the Schwarzian. The difference between the two casts comes from the initial conditions. Those for $P$ are dictated by the behaviour at $U=$-$\infty$ but those for $Q$ involve also $S$ in (II.10).
V Illustration by the Pöschl - Teller profile
We illustrate our general theory by the Pöschl - Teller profile PTeller ; Plyushchay09 ; ChaKar ; DM-1 in $D=1$,
${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{m(m+1)}{\cosh^{2}U}\,.$ (V.1)
The Sturm-Liouville equation (II.6) is solved by Legendre functions. DM trajectories are obtained when $m$ is an integer and we get Legendre polynomials,
$P(U)=(-1)^{m}P_{m}(\tanh U)\,.$ (V.2)
The general geodesic, (III.6), combines the DM solution $P(U)$ and the non-polynomial $Q(U)$ in (III.3), shown FIG.1.
Figure 1: The solutions of the Sturm-Liouville eqn. (II.6) for Pöschl - Teller , shown for DM parameters. With initial conditions (II.19) we get Legendre polynomials, $P(U)=P_{m}(\tanh U)$. The second, independent solution, $Q(U)=Q_{m}(\tanh U)$, is a Legendre function which does not satisfy the initial condition (II.19).
Turning to BJR cf. (II.7), the Souriau matrix $S$ in (II.10) is regular between two subsequent zeros of $P$ but diverges at the junction points. For $m=1$, for example,
$S_{m=1}(u)=u-\coth u\,$ (V.3)
obtained for the choice $u^{\pm}_{0}\approx\pm 1.2$, which is regular both in $I_{-}=\big\{u<0\big)\}$ and in $I_{+}=\big\{u>0\big)\}$, but diverges at $u=0$, as depicted in FIG.2a. Similarly for $m=2$ shown in FIG.2b,
$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$ (V.4)
is singular where the denominator vanishes.
Figure 2: The Sturm-Liouville solutions $\bf P$ of (II.6) shown with wave numbers $\bf m=1$} and $\bf m=2$} for Pöschl - Teller in $D=1$. They yield also DM trajectories written in Brinkmann coordinates. The Souriau matrix S diverges where $\bf P=0$, but the pull-back to Brinkmann $\color[rgb]{0,0.5,0}\bf Q=PS$} is regular for all $U$.
$Q_{m}=P_{m}S_{m}$ in (III.3), shown in FIG.2 in green, is a second solution of the Sturm-Liouville equation (II.6). It is regular for all $U$, vanishes where the Souriau matrix does, and exhibits VM behaviour outside the Wave zone. $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$. For $m=1$ and $2$ we have, for example,
[…]
The $P(U)$ found above provide us with DM geodesics which are consistent with (II.18). The zeros of $P$ are caustic points as it is manifest in FIG.3.
Figure 3: For a DM wave in $D=1$ transverse dimension Brinkmann trajectories focus to the caustic points determined by $P_{m}(U_{k})=0$, $k=1,\dots,m$. For $m=1$ there is one focal point, at the origin, and for $m=2$, we have two of them, determined by $\tanh U_{\pm}=\pm{1}/{\sqrt{3}}$.
Turning to symmetries, the coefficients of the globally defined Carroll generators $P$ and $Q$ in (III.1), listed in (V.6) and (V) are depicted in FIGs. 4 and 5.
$\Theta_{m=1}^{Brink}$ $=$ $h\frac{\partial}{\partial V}+c\Big({\tanh U}\,\frac{\partial}{\partial X}-(\operatorname{sech}^{2}U)X\frac{\partial}{\partial V}\Big)$ (V.6)
$+$ $b\Big({\big(U\tanh U-1\big)}\frac{\partial}{\partial X}-\big(\tanh U+U\operatorname{sech}^{2}U\big)X\frac{\partial}{\partial V}\Big)\,.$
$\Theta_{m=2}^{Brink}\!$ $=$ $\!h\frac{\partial}{\partial V}+c\left(\big(3\tanh^{2}U-1\big)\frac{\partial}{\partial X}-\big(6\tanh U\operatorname{sech}^{2}(U)\big)X\frac{\partial}{\partial V}\right)$
$\thinspace{+}b\left(\frac{1}{4}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\frac{\partial}{\partial X}-\frac{1}{2}\big(1-3\operatorname{sech}^{2}U+3U\operatorname{sech}^{2}U\tanh U\big)X\frac{\partial}{\partial V}\right).\qquad$
(a)                        (b)
Figure 4: For wave number $\bf m=1$}, the (a) translation generator ${\bf P}_{1}$ [which generates also a trajectory] is odd. It has one zero, at $U=0$. (b) the generator of a Carroll boost, ${\bf Q}_{1}={\bf P_{1}S_{1}}$, is even and has two, symmetrically positioned zeros.
(a)                        (b)
Figure 5: Infinitesimal Carroll generators (V) with wave number $\bf m=2$}. (a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$. (b) For a boost, ${\bf Q}_{2}={\bf P_{2}S_{2}}$ is odd and has 33 zeros.
VI Graphic representation of the symmetry generators
The symmetry-generating vectors along the trajectories shown in FIGs.4 and 5 are conveniently viewed in a co-moving tangent space – which is a plane carried along the trajectory. In FIGs. 69 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of the symmetry generators $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$ in (III.1), spelt out in (V.6)-(V).
Figure 6: Infinitesimal translations $({\bm{\xi}},\eta)$ along a DM Brinkmann trajectory ${\bm{X}}(U)$ with $\bf m=1$}, shown in the co-moving tangent space. The trajectory is “hidden” in the yellow blob.
FIG.6 shows that the translation vector starts at $\bf U=-\infty$} from $\color[rgb]{.75,.5,.25}{\bm{\xi}}=-1,\eta=0$}. It leaves the caustic point for $\bf U=0$} at ${\bm{X}}=0,\,V=0$ invariant. Arriving into the Afterzone, the translation vectors change sign and end, for $U=+\infty$, at $\color[rgb]{0,1,1}{\bm{\xi}}=+1,{\eta}=0$}. Thus they act again as usual translations – but with reversed sign.
Figure 7: Infinitesimal boosts along along the trajectory “hidden behind the yellow blob” with $\bf m=1$}, shown in the co-moving tangent space.
In FIG.7, the longest arrow (in brown) is both the initial and the final boost vector for $\bf U=\mp\infty$}. The shortest red arrow $({\bm{\xi}}=1,{\eta}=0)$ is reached for $\bf U=0$} at the caustic point ${\bm{X}}=0,V=0$, where the boost shifts all trajectories by the same amount.
Figure 8: Translations act along the $\bf m=2$} DM trajectory, see (V). The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
According to FIG.8, the translation generator (with $c=1/2$) is, for $U=$-$\infty$, $({\bm{\xi}}=1,\,\eta=0)$. FIGs. 4 and 5 confirm that, consistently with (V.6), the two focal points of the trajectory, at $U_{\mp}$, are left invariant. For $U=+\infty$ the “8-shaped” curve returns to $({\bm{\xi}}=1,\,\eta=0)$, where it started from.
Figure 9: Infinitesimal Carroll boosts in Brinkmann coordinates along a chosen DM trajectory marked by a yellow blob, with wave number $\bf m=2$}.
Boosts along the $m=2$ trajectory are shown in FIG.9. The longest vector on the left (in brown) is, consistently with (V.6) and (V) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$. The two shortest blue arrows (${\bm{\xi}}=\pm 1,{\eta}=0$) show the boost acting as a translations by $\mp b$ at the caustic points $U_{\mp}$. Combining FIGs.4 with FIG.6 provides us with the outfolding to 3 dimensions, shown in FIG.10.
Figure 10: Translations implemented along the trajectories with wave numbers $m=1$ and $m=2$.
The translations implemented along the trajectories vanish where $P(U)=0$. They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
Figure 11: Carroll boosts implemented along the trajectories with wave number $m=1$ and $m=2$.
Carroll boosts implemented along the trajectories behave as pure translations at P(U)=0P(U)~=~0. They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
Conversely, we duly recover the Pöschl - Teller profile (V.1) from the Sturm-Liouville solutions by using the Schwarzian formulas in sect.IV. We show this here for the simplest, $m=1$ solution (V.5), with the cast
$\varphi_{1}(U)=P(U)=-\tanh U,\qquad\varphi_{2}(U)=Q(U)=1-U\tanh U\,.$ (VI.1)
The Wronskian equals to (1)(-1) everywhere. Then inserting $f^{\prime}(U)=-\cosh^{2}U+1\,$ into the Schwarzian (IV.4) reproduces the $m=1$ profile ${\mathcal{A}}(U)=-\frac{8}{\cosh^{2}(U)}$ in (V.1).
VII Conclusion
Our results shed further light on the intimate relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.19) yields our geodesics. The general expression (III.6) which involves both DM and VM is one of our main results.
When the parameter $m$ is an integer, then the wavezone accommodates $m$ half-waves and we get DM. For $m=2\ell+1$ $P(U)$ is odd and we do get net displacement. For $m=2\ell$, though, $P(U)$ is even and the particle returns, after non-trivial motion in the wave zone, to its initial position : we get DM with no final displacement, consistently with (II.18) and confirmed by FIGs.3, 4a and 5a. The Brinkmann trajectories are focused at the zeros of $\det(P)$.
DM trajectories have zero momentum and zero non-relativistic energy DM-1 . They can be found by putting ${\bm{p}}_{0}=0$ into (III.6). Turning on the momentum, ${\bm{p}}_{0}\neq 0$, converts the trajectory to VM.
Both BJR and Brinkmann coordinates have their advantages and drawbacks.
The merit of BJR is simplicity Sou73 ; Carroll4GW . Our initial conditions (II.19) imply ${\bm{p}}=0$, and the “motion” reduces merely to a fixed point determined by the conserved boost momentum ${\bf k}={\bm{x}}_{0}$ and $v_{0}$ — the initial position (II.13)-(II.17). The dynamics is carried by $P(u)$. The price to pay is that the BJR description is valid only between the mandatory zeros of $\det(P)$, where the $S$-matrix diverges Sou73 ; LongMemory ; EZHRev , and solutions in the adjacent domains have to be fitted together EZHRev .
The advantage of the Brinkmann form (II.3)–(II.5) is that it replaces the singular Souriau matrix $S$ in (II.10) by the globally defined Sturm-Liouville solutions $P$ and $Q=PS$, (II.6) and (III.3), respectively. For zero momentum, (II.16), the non-DM solution $Q$ is turned off, leaving us with $P$ alone which plays its second, trajectory-role. The drawback is that finding those solutions may be laborious QLdPT ; DM-1 ; Sila-PLB . In conclusion, the main difficulty boils down, in both coordinate systems, to solving a Sturm-Liouville equation.
Switching to Brinkmann coordinates allowed us to extend the locally given Carrollian Killing vectors Sou73 to the whole spacetime. These vectors are directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by combining the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.6), highlighting the intimate relationship of geodesics with Carroll symmetry EZHRev ; GenRelGrav .
The Brinkmann and BJR forms have also been compared by Harte Harte from a different point of view. Showing that the BJR plane wave metric is a generalisation of perturbative weak gravitational waves, he discussed memory effects in a perturbative manner.
Figs. 611 illustrate the curious behavior of the Carroll generators along the trajectories.
Further insight is provided by the Schwarzian framework Ovsienko ; ZZH ; Vlahakis . This approach confirms that $Q=PS$ is indeed a globally defined, 2nd solution of the Sturm-Liouville problem which does not involve the singular Souriau matrix $S$.
Our results go beyond the toy dimension $D=1$ dimensional profiles and the Pöschl - Teller example. Similar results hold for the Scarf profile in $D=2$, which provides us with an analytic approximation of the physically relevant flyby case ZelPol ; Sila-PLB .
Gravitational waves and their observable effects are also formulated by means of Bondi metric Winicour which relates the gravitational emission and the mass loss. The Bondi metric is axially symmetric and asymptotically flat. In general, our plane waves are not asymptotically flat, though. However, both metrics share similar properties like having a covariantly constant null Killing vector which determines the direction of wave propagation. The transverse part of the Bondi metric is quadratic in the radial coordinate which defines the luminosity distance. In Harte , the BJR plane wave metric is used to compute similar physical properties. To clarify the differences and the similarities between the Bondi formalism with ours here requires more effort and will be reported elsewhere.
Supersymmetry MSUSY sheds further light on the Memory Effect. Further details will be presented in a comprehensive review ZEHPR .
Acknowledgements.
The authors are indebted to J. Balog, J. Ben Achour and Z. Silagadze for correspondance. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320). ME is supported by Istanbul Bilgi University research fund (BAP) with grant no: 2024.01.009 and by The Scientific and Technological Research Council of Turkey (TÜBITAK) under grant number 125F021. PMZ and PAH benefitted of TÜBITAK 2221 grant and hospitality at Bilgi University during their visit in Istanbul.